OpenStudy (anonymous):

(HELP, Medal for help) 2x/3b X 9bx/4x^3

OpenStudy (anonymous):

Multiply across Then you get: (18b(x^2)) / (12b(x^3) Simplify: (3) / (2x)

OpenStudy (anonymous):

\[\frac{ 2x }{ 3b }\times \frac{ 9bx }{ 4x^3 }\]

OpenStudy (anonymous):

Is it like that?

OpenStudy (anonymous):

Btw @ivy_green is wrong. You don't cross multiply...it's not 2x/3b = 9bx/4x^3...

OpenStudy (anonymous):

Well thanks for looking at least looking at it and telling me about the cross multiple guys.

OpenStudy (anonymous):

Not to cross multiply just multiply across (2x)(9bx) / (3b)(4x^3)

OpenStudy (anonymous):

Just multiply as you would a normal fraction...no cross multiplication as I explained earlier.